OpenStudy (anonymous):

Will fan AND GIVE MEDAL!!! There are 10 boys and 8 girls on the cross country team the ratio of boys to girls participating in track and field is proportional to the ratio of boys to girls on the cross country team. Which could be the number of boys to girls participating in track and field? A. 12 boys and 10 girls B. 14 boys and 12 girls C 15 boys and 12 girls D. 16 boys and 14 girls

OpenStudy (anonymous):

@Data_LG2 Can you explain this maybe

OpenStudy (anonymous):

It looks like all except c are sort of correct... I'm not much help, I'm afraid...

OpenStudy (anonymous):

This is the same process like you did before. #boys in a country team : #girls in a country team = #boys in t&f : # girls in t&f you have: \(\sf \large \frac{10\ boys \div 2}{8 \ girls \div 2} = \frac{5 \ boys}{4 \ girls}\) now which of the following is equal to 5/4 (boys/girls) ?

OpenStudy (anonymous):

I mean which of the choices**

OpenStudy (anonymous):

5/4 = 1.25

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

So confused :C

OpenStudy (anonymous):

the total = 10+8=18 boys = 10/18=55.56 girls = 8/18=44.44

OpenStudy (anonymous):

yes, that's right.. which of your choices is equal to 5/4 or 1.25 , if you'll calculate for their ratio? e.g. choice A: 12 boys and 10 girls = 12/10 = 6/5 which is not equal to 5/4 so this is not the answer. try doing this with the rest of your choices and tell me which one yields 5/4

OpenStudy (anonymous):

c is the answer

OpenStudy (anonymous):

15+12=27 boys = 15/27= 55.56

OpenStudy (anonymous):

Ya C is answer I just divided it. THANK U guys!! @AJ01 @Data_LG2 ill fan and vmedal one

OpenStudy (anonymous):

YW